# Orbital Description of Landau Levels

### v2

@@ Abstract
where $k_{\pm}=k_{x}\pm ik_{y}$. When $\Delta=-v_{F}^{2}/2(\alpha+\beta)$, both bands satisfy the trace and determinant conditions at any $\mathbf{k}$ with $g_{\mu\nu}(\mathbf{k})=(1/2)\left|\mathcal{B}(\mathbf{k})\right|\delta_{\mu\nu}$. Such an ideal local band inversion always lead to local Chern number $\mathcal{C}_{\text{local}}=1$ and local average trace condition $\mathbb{T}_{\text{local}}=1$ when the integration is applied around the band inversion point. This well explains that the orbitals $s$, $p_{-}$, $p_{+}$ constitute a complete subspace to describe the LLL and 1LL shown in Fig. 1 and Fig. 3. For the LCB, there is only an inversion of the band between $s$ and $p_{-}$ at $\Gamma$, thus $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ concentrate around $\Gamma$, and the integral of them around $\Gamma$ gives the Chern number of the band $\mathcal{C}=1$ and $\mathbb{T}\approx 1$. For the 1CB, there are $p_{-}$ and $s$ band inversion at $\Gamma$, as well as $p_{-}$ and $p_{+}$ band inversion at $K$, the integral of $\mathcal{B}(\mathbf{k})$ around $\Gamma$ and $K$ is $-1$ and $+1$, respectively, while local $\mathbb{T}$ around $\Gamma$ and $K$ are both nearly 1, then the sum of them finally leads to the band Chern number $\mathcal{C}=2-1=1$ and $\mathbb{T}$ = $\left|2\right|+\left|-1\right|$ = 3. Such processes could be applied to higher Chern bands. As shown in Fig. 1(b,e), the band inversion happened at $M$ between $p_{+}$ and higher $d$ orbital provide local Chern number $\mathcal{C}_{\text{local}}=1$ and local $\mathbb{T}_{\text{local}}=1$ at each $M$, combined with the $p_{+}$ and $p_{-}$ band inversion at $K$, we get Chern number $\mathcal{C}=3-2=1$ and $\mathbb{T}$ = $\left|3\right|+\left|-2\right|$ = 5 for second Chern bands (namely, the third band counting from below), which is the ideal number for second LL. The quantum geometry of these constructed Chern bands matches with that of the generalized LL [76]. It is worth mentioning that the Chern bands from such construction only have perfect quantum weight at the penalty of Berry curvature flatness, since $\mathcal{B}(\mathbf{k})$ always concentrates at the local band inversion point. To obtain the FCI state in the partially filled Chern bands, a delicate balance between $\mathcal{B}(\mathbf{k})$ flatness and ideal $\mathbb{T}$ should be considered. We also calculate the minimal model with short-range hopping up to third nearest neighbors, where similar band structure and quantum geometry are obtained, and the many-body topological states retain [53]. The construction of the projected Hamiltonian within the triangular lattice here can be directly generalized to the square lattice, provided that $\mathbf{S}(\mathbf{r})$ exhibits $C_{4}$ symmetry. Furthermore, our results could be applied to cold atom systems [77], since only short-range density-density interactions are required.
@@
We point out that the orbital construction of LL here is consistent with the higher vortexability for 1LL proposed in Ref. [34]. The 1CB here is not self-vortexable because of its large value of $\mathbb{T}$. When we combine LCB and 1CB together, we can see a nearly ideal $\mathbb{T}=2.2$ for the lowest two Chern bands, since they originate from band inversion between $p_{-}$ and $p_{+}$ at $K$ point.
@@ Acknowledgements.
Acknowledgment. We thank Biao Lian and Dung-Hai Lee for helpful discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600 and No. 2019SHZDZX01. H.W. and R.S. contributed equally to this work.

### v3

@@ Abstract
where $k_{\pm}=k_{x}\pm ik_{y}$. When $\Delta=-v_{F}^{2}/2(\alpha+\beta)$, both bands satisfy the trace and determinant conditions at any $\mathbf{k}$ with the quantum metric given by $g_{\mu\nu}(\mathbf{k})=(1/2)\left|\mathcal{B}(\mathbf{k})\right|\delta_{\mu\nu}$. Such an ideal local band inversion consistently yields a local Chern number $\mathcal{C}_{\text{local}}=1$ and a local average trace condition $\mathbb{T}_{\text{local}}=1$ when integrated over the region surrounding the inversion point. This framework explains why the orbitals $s$, $_{-}, p+p_{+}$ form a complete basis for describing the LLL and 1LL, as illustrated in Fig. 1 and Fig. 3. For the LCB, a single band inversion occurs between the $s$ and $p_{-}$ orbitals at the $\Gamma$ point. As a result, both $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ are concentrated around $\Gamma$, and the integral of them around $\Gamma$ yield a Chern number $\mathcal{C}=1$ and an average trace condition $\mathbb{T}\approx 1$. For the 1CB, two band inversions occur: one between $p_{-}$ and $s$ at $\Gamma$, and the other between $p_{-}$ and $p_{+}$ at $K$. The integral of $\mathcal{B}(\mathbf{k})$ around $\Gamma$ gives 1-1, and around $K$, it gives +1+1; the corresponding local $\mathbb{T}$ are both close to $1$. This results in a net Chern number $\mathcal{C}=2-1=1$, and a total trace condition $\mathbb{T}=\left|2\right|+\left|-1\right|=3$. This construction can be extended to higher Chern bands. As shown in Fig. 1(b,e), the band inversion at $M$ between $p_{+}$ and upper $d$ orbital contributes a local Chern number $\mathcal{C}_{\text{local}}=1$ and $\mathbb{T}_{\text{local}}=1$ at each $M$. Combined with the band inversion between $p_{+}$ and $p_{-}$ at $K$, this yields a total Chern number $\mathcal{C}=3-2=1$ and $\mathbb{T}=\left|3\right|+\left|-2\right|=5$ for the second Chern bands (i.e., the third band from the bottom), which matches the ideal value for the second LL. These results show that the quantum geometry of the constructed Chern bands closely matches that of generalized LL [73]. It is worth mentioning that the constructed Chern bands only have a perfect quantum weight at the penalty of Berry curvature flatness, since $\mathcal{B}(\mathbf{k})$ always concentrates at the local band inversion point. To obtain the FCI state in the partially filled Chern bands within exact diagonalization, a delicate balance between $\mathcal{B}(\mathbf{k})$ flatness and ideal $\mathbb{T}$ should be considered.
We also examine a minimal model that incorporates short-range hopping up to third-nearest neighbors, yielding a similar band structure and quantum geometry. Importantly, the many-body topological states remain robust in this setting [59], making it suitable to implementation in cold atom systems [82, 83, 84, 85], since only short-range density-density interactions are required. Optical lattices with synthetic gauge fields offer a controllable platform for realizing such orbital-based models with tunable interactions. Crucially, the lattice formulation is not a mere simplification but a central advantage, opening a practical route toward experimental realizations of fractional Chern insulators and non-Abelian phases in ultracold gases.
@@
We note that the orbital construction of LL presented here is consistent with the concept of higher vortexability for the 1LL, as proposed in [40]. Specifically, the 1CB in our model is not self-vortexable, due to its large quantum metric trace $\mathbb{T}$. However, when the LCB and 1CB are considered together, we observe a nearly ideal combined trace $\mathbb{T}=2.2$ for the two lowest bands. This behavior originates from the band inversion between the $p_{-}$ and $p_{+}$ at the $K$ point. Our results thus offer a new perspective on understanding the vortexability of the lowest few LLs from a lattice orbital viewpoint.
Our orbital-based construction of LLs provides a complementary approach to QH lattice models. The well-known Kapit–Mueller (KM) model offers an exact lattice realization of LLL wavefunctions through engineered hopping terms [86, 87]. In contrast, our scheme relies on sequential band inversions among local Wannier orbitals of $s$, $_{-}, p+p_{+}$ character. This mechanism not only reproduces the LLL but also generates higher LLs within a unified multiband framework, thereby offering direct access to non-Abelian states inherent to higher levels.
@@ Acknowledgements.
Acknowledgments—We thank Biao Lian and Dung-Hai Lee for helpful discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission. H.W. and R.S. contributed equally to this work.